# with Language Model Diagnostic Assistance

### v1

@@ Abstract
In the emergency department (ED), patients undergo triage and multiple laboratory tests before diagnosis. This process is time-consuming, and causes ED crowding which significantly impacts patient mortality, medical errors, staff burnout, etc. This work proposes (time) cost-effective diagnostic assistance that explores the potential of artificial intelligence (AI) systems in assisting ED clinicians to make time-efficient and accurate diagnoses. Using publicly available patient data, we collaborate with ED clinicians to curate MIMIC-ED-Assist, a benchmark that measures the ability of AI systems in suggesting laboratory tests that minimize ED wait times, while correctly predicting critical outcomes such as death. We develop ED-Copilot which sequentially suggests patient-specific laboratory tests and makes diagnostic predictions. ED-Copilot uses a pre-trained bio-medical language model to encode patient information and reinforcement learning to minimize ED wait time and maximize prediction accuracy of critical outcomes. On MIMIC-ED-Assist, ED-Copilot improves prediction accuracy over baselines while halving average wait time from four hours to two hours. Ablation studies demonstrate the importance of model scale and use of a bio-medical language model. Further analyses reveal the necessity of personalized laboratory test suggestions for diagnosing patients with severe cases, as well as the potential of ED-Copilot in providing ED clinicians with informative laboratory test recommendations. Our code is available at https://github.com/cxcscmu/ED-Copilot.
@@ 1 Introduction
Emergency Department (ED) crowding represents a critical challenge in healthcare, significantly impacting morbidity, mortality, medical error, staff burnout, and incurring excessive costs (Sartini et al., 2022). Despite the documented effects of ED crowding, this issue remains underappreciated and inadequately addressed in healthcare systems. As a sentinel indicator of health system functioning, the ED is vital for providing timely access to care for patients who are seriously ill or injured (Savioli et al., 2022).
@@
One key area to address ED crowding, as identified by the American College of Emergency Physicians, is to enhance throughput—the efficacy and efficiency of care delivery in the ED (Jarvis, 2016; DeAnda, 2018). A crucial factor affecting the throughput is long ED wait-times due to the laboratory testing process (Li et al., 2015). Patients frequently experience lengthy waits before their tests are ordered, followed by additional hours waiting for the test results. Clinicians can only make a diagnosis and decide on next steps for treatment after recieving these results. Furthermore, studies have indicated that approximately 40 to 60% of laboratory tests conducted in the ED are unnecessary (Miyakis et al., 2006), resulting in increased wait-times and reduced throuhput.
@@
This paper proposes an artificial intelligence (AI) “Co-Pilot” system intended to offer (time) cost-effective diagnostic assistance in the ED. The system minimizes ED length of stay (LOS) – a proxy for wait-times – by automatically suggesting laboratory tests after patient triage. Minimizing wait-time via ED LOS serves to improve throughput, and hence crowding. Additionally, the system identifies patients at high risk of deterioration who require rapid intervention. By selecting informative laboratory tests, the system streamlines the diagnostic process, reducing time-cost (i.e., LOS) while ensuring excellent outcomes in terms of identifying high-risk patients.
@@
To enable the machine learning (ML) community to study (time) cost-effective diagnostic assistance, we collaborate with ED clinicians and utilize publicly available patient records in MIMIC-IV (Johnson et al., 2023b) and other ED benchmarks (Xie et al., 2022) derived from MIMIC-IV, to develop MIMIC-ED-Assist, a benchmark that tests the ability of AI systems to help ED practitioners navigate the diagnostic space. Specifically, our ED clinicians identified two prediction targets: critical outcome, which includes patient death and ICU transfer (Levin et al., 2018), and lengthened ED stay, defined as ED LOS exceeding $24$ hours. Both reflect patient risk. Accurately identifying patients with these outcomes reduces time-cost by allowing clinicians to intervene in a timely fashion, and efficiently allocate resources. Furthermore, to mirror real-world ED practices, our clinicians grouped individual laboratory tests into commonly performed ED laboratory groups, e.g. complete blood count (CBC). MIMIC-ED-Assist then tests AI systems on their ability to recommend the most informative groups to make accurate diagnostic suggestions while minimizing the total time required to perform these tests, hence reducing LOS.
@@
With MIMIC-ED-Assist, we develop ED-Copilot, which learns to suggest a series of laboratory groups, with minimum total time cost, and flags patients at high risk of critical outcome or lengthened ED stay. ED-Copilot first linearizes (i.e., converts to text) patient information, including demographic, triage, and laboratory test results into a text sequence. It then fine-tunes a bio-medical pre-trained language model BioGPT (Luo et al., 2022) to suggest future groups and predict the two outcomes. Next, we use a reinforcement learning (RL) framework introduced by Yu et al. (2023) to teach BioGPT to dynamically recommend the subsequent, most informative laboratory group based on prior lab and triage information. ED-Copilot effectively flags high-risk patients while minimizing the number of laboratory groups to reduce ED LOS, our proxy for wait-time.
@@
Experiments on MIMIC-ED-Assist show that ED-Copilot is not only more accurate than state-of-the-art tree models trained on laboratory test results for both outcomes, but does so while using significantly fewer tests. ED-Copilot halves time cost of laboratory testing from four hours to two hours. Minimizing number of lab tests also has the benefit of reducing financial cost. We also perform ablation studies to confirm the benefits of our feature linearization technique and the bio-medical pre-trained language model backbone.
Our analyses confirmed the necessity of a personalized cost-effective diagnostic assistant. Non-personalized cost-effective methods minimize time-cost by selecting laboratory groups that maximize overall predictability. However, patients present differently, and require personalized care, particularly if they are at higher risk of critical outcome. By personalizing lab group suggestions, ED-Copilot provides more accurate diagnostic assistance for patients with unique needs. Furthermore, we show that ED-Copilot makes reasonable lab group suggestions despite results only available for tests the patient actually receives, due to the limitation of MIMIC-ED-Assist being a retrospective offline benchmark.
The main contributions of this paper are as follows.
  1. 1.
    We work with ED clinicians to develop MIMIC-ED-Assist, a publicly accessible benchmark designed to advance research in ED diagnostic assistance.
  2. 2.
    We develop ED-Copilot, an RL-trained language model that enhances ED workflow by suggesting informative laboratory groups and flagging high-risk patients using a pre-trained language model backbone BioGPT.
  3. 3.
    A myriad of experiments and analyses demonstrate that ED-Copilot improves accuracy over state-of-the-art classifier while reducing ED wait-time by 50%.
@@ 2 Related Work
Healthcare benchmarks. Researchers have spent considerable effort in converting raw electronic health records (EHRs) into large-scale open-source datasets to ensure easy access to high-quality medical data. A notable example is the Medical Information Mart for Intensive Care (MIMIC) database (Johnson et al., 2023b) which provides patient information such as measurements, laboratory orders, and treatments, ranging from the ED to inpatient care, including the intensive care unit (ICU). MIMIC has led to the development of a range of related prediction benchmarks and models (Harutyunyan et al., 2019; Purushotham et al., 2018; Wang et al., 2020) focused on the ICU. Xie et al. (2022) took a step towards filling this gap by using the MIMIC-IV-ED (Johnson et al., 2023a) database to build a ED-focused benchmark. Their dataset includes ED triage information, and various clinical outcomes such as hospitalization that interest clinicians, and impact ED LOS.
AI models for healthcare. Alongside developing benchmarks, the ML community has also tried to build models to predict various clinical outcomes. Traditional ML methods (e.g., random forests, gradient boosting, and their variants (Breiman, 2001; Chen & Guestrin, 2016; Agarwal et al., 2022, 2023)), along with deep learning (DL) have been used to predict pneumonia (Kang et al., 2020), and septic shock in the ICU (Wardi et al., 2021). Other works use interpretable models to provide diagnostic assistance in the ED such as identifying traumatic brain injury (Tan et al., 2022; Kornblith et al., 2022).
A line of work closely related to our paper focuses not only on prediction but also on AI for cost-effective medicine. For instance, Bejnordi et al. (2017) showed DL leads to faster analysis of pathology laboratory results; Komorowski et al. (2018) proposed an “AI clinician” to learn optimal dosing strategies for treating sepsis; Yu et al. (2023) focused on minimizing financial costs associated with laboratory testing while maximizing prediction accuracy. Specifically, Yu et al. (2023) used RL to sequentially select laboratory groups based on a patient’s observed test results to optimize this (financial) cost-accuracy trade-off. They validated the accuracy and cost-effectiveness of their approach on multiple clinical tasks such as diagnosing kidney injury. Researchers have also begun to explore the use of large language models (LLMs) in medical applications. LLMs have been used to extract clinical concepts (Luo et al., 2022; Yang et al., 2022, 2021), and facilitate medical question answering (Yagnik et al., 2024; Singhal et al., 2023). However, the integration of LLMs into ED workflows, particularly in addressing ED crowding, remains unexplored.
@@ 3 MIMIC-ED-Assist Benchmark
We discuss how we curate MIMIC-ED-Assist in collaboration with clinicians, a benchmark with comprehensive laboratory information that allows for the development of AI systems that can tackle long ED wait-times via minimizing laboratory testing time-cost.
@@
Task description. We focus on the following two prediction tasks. 1) Flagging patients at high risk of critical outcome (i.e., death and ICU transfer), and lengthened ED. 2) Providing time cost-effective diagnostic assistance by recommending the next most medically informative laboratory group while simultaneously minimizing the time-cost of these groups. As discussed in detail below, both tasks help clinicians optimize their workflow, and serve to achieve the larger goal of reducing ED LOS.
@@
Data pre-processing. We filter out non-hospitalized ED patients from MIMIC-IV-ED, as laboratory results are only available for admitted patients. We only focus on adults and remove patients younger than $18$ years old. Further, we exclude patients with missing triage information. This step is necessary since clinicians order laboratory groups depending on triage information. We simplify our task by removing patients who receive the same test multiple times, approximately $1.5\%$ of all patients.
@@
Clinically relevant outcomes. We construct two prediction outcomes that help optimize ED workflow and reduce ED LOS. (1) Critical condition, which refers to death during hospitalization or transfer to an ICU within 12 hours. (2) Lengthened ED stay, indicating if ED LOS exceeding $24$ hours. These outcomes were chosen as prediction targets in close collaboration with ED clinicians for the following reasons. If a patient is identified to be at high-risk of a critical outcome, clinicians will prioritize treatment and resources for this patient. Lengthened ED stay is typically correlated with the complexity of a patient’s case. Flagging patients at high risk of lengthened ED stay can enable timely intervention, and reduce ED LOS. While these two outcomes are correlated, they cover different aspects of patient care. For example, patients at high risk of critical outcomes often should be hospitalized quickly, while patients with complications do not necessarily have severe cases. As such, healthcare providers often require different diagnostics and resource management for each of these two tasks.
@@
Triage feature selection. We select a number of triage features in close collaboration with ED clinicians. Triage features are available for every patient before any laboratory tests are ordered. Specifically, our ED clinicians chose $9$ triage variables that are available at the beginning of the patient encounter. These variables included patient demographics, medical history, vital signs, and chief complaints from triage (i.e., natural language description of symptoms).
@@
Laboratory test selection. We only include laboratory tests performed in the ED. For simplicity, we exclude laboratory tests received by less than $5\%$ of patients and leave further examination of rare laboratory tests in future research. This process results in a total of $68$ available laboratory tests. While there are $68$ laboratory tests, ED clinicians rarely order individual tests for a patient. Rather, they order groups or groups of tests based on a patient’s signs, symptoms, and risk factors. For example, a clinician might order a basic metabolic group or a complete blood count. To reflect this clinical practice, we work with ED clinicians to categorize these 68 tests into 12 distinct groups, representing common groupings in clinical settings. See Appendix A for all laboratory tests, and their assigned groupings. On average each patient has $4.7$ laboratory groups. Consequently, our dataset contains numerous missing values, each representing a laboratory test not administered to a specific patient.
@@
ED LOS. Our benchmark records when each group is ordered, and assigns its average completion time as its ’time-cost’. ED LOS depends on the time-costs of the administered groups and can be modeled in different ways. For example, sequential tests result in ED LOS being the sum of time costs, whereas parallel tests imply ED LOS is equal to the group with the largest time-cost. MIMIC-ED-Assist does not specify how to approximate ED LOS from these time-costs, and instead provides this flexibility to researchers and practitioners.
@@
Table 1: Statistics of MIMIC-ED-Assist. It includes information from patient triage and lab tests. Lab tests are grouped into 12 groups by ED clinicians based on how they are commonly ordered. An ED visit has a critical outcome if the patient is transferred to ICU or there is an inpatient mortality. ED stay is lengthened if the length of stay (LOS) exceeds 24 hours.
@@
##\## of lab variables 67
##\## of lab groups 12
Avg ##\## of lab groups per patient 4.7
@@
Data availability. Our pipeline to curate MIMIC-ED-Assist from the MIMIC-IV datasets will be made available. MIMIC-IV is publicly available, and then one can use our pipeline to create MIMIC-ED-Assist.
Limitations. There are several limitations of MIMIC-ED-Assist, mainly due to data availability challenges in public patient data (1) MIMIC-ED-Assist is derived from MIMIC-IV which only has laboratory results for hospitalized ED patients. Thus, our dataset is not reflective of the entire population that visits the ED, but instead is biased towards those with more severe issues (hospitalized). (2) As an offline benchmark derived from past patient records, all data is retrospective. When training models to recommend future laboratory groups, one can only use laboratory tests patients actually received, otherwise no testing results are available. This leads to common challenges of offline benchmarks and may result in models learning sub-optimal laboratory group assignment, and unobserved confounding. (3) MIMIC-IV is collected at the Beth Israel Deaconess Medical Center, and may not reflect the distribution of other healthcare systems. Caution is needed when generalizing models and insights derived from MIMIC-IV to other places.

### v2

@@ Abstract
In the emergency department (ED), patients undergo triage and multiple laboratory tests before diagnosis. This time-consuming process causes ED crowding which impacts patient mortality, medical errors, staff burnout, etc. This work proposes (time) cost-effective diagnostic assistance that leverages artificial intelligence systems to help ED clinicians make efficient and accurate diagnoses. In collaboration with ED clinicians, we use public patient data to curate MIMIC-ED-Assist, a benchmark for AI systems to suggest laboratory tests that minimize wait time while accurately predicting critical outcomes such as death. With MIMIC-ED-Assist, we develop ED-Copilot which sequentially suggests patient-specific laboratory tests and makes diagnostic predictions. ED-Copilot employs a pre-trained bio-medical language model to encode patient information and uses reinforcement learning to minimize ED wait time and maximize prediction accuracy. On MIMIC-ED-Assist, ED-Copilot improves prediction accuracy over baselines while halving average wait time from four hours to two hours. ED-Copilot can also effectively personalize treatment recommendations based on patient severity, further highlighting its potential as a diagnostic assistant. Since MIMIC-ED-Assist is a retrospective benchmark, ED-Copilot is restricted to recommend only observed tests. We show ED-Copilot achieves competitive performance without this restriction as the maximum allowed time increases. Our code is available at https://github.com/cxcscmu/ED-Copilot.
@@ 1 Introduction
Emergency Department (ED) crowding represents a critical challenge in healthcare, significantly impacting morbidity, mortality, medical error, staff burnout, and incurring excessive costs (Sartini et al., 2022). Despite the documented effects of ED crowding, this issue remains inadequately addressed in healthcare systems. An efficient and effective ED is vital for providing timely care to severely ill or injured patients (Savioli et al., 2022).
@@
One key area to address ED crowding, as identified by the American College of Emergency Physicians, is to enhance throughput—the efficacy and efficiency of care delivery in the ED (Jarvis, 2016; DeAnda, 2018). A crucial factor affecting throughput is the laboratory testing process, where patients often face lengthy waits for tests to be ordered and completed, delaying diagnoses and treatment decisions  (Li et al., 2015). Studies also show that 40 to 60% of ED laboratory tests are unnecessary (Miyakis et al., 2006), further exacerbating wait times.
@@
This paper proposes an artificial intelligence “Co-Pilot” system intended to offer (time) cost-effective diagnostic assistance in the ED. This system should aid diagnosis and minimize ED length of stay (LOS), i.e., wait times, by suggesting laboratory tests after patient triage. Further, it should help with resource management and planning by identifying severely ill patients who require rapid intervention. That is, by selecting informative tests, the system streamlines the diagnostic process, reducing LOS while improving outcomes, particularly for high-risk patients.
@@
To support the machine learning (ML) community in developing a time-cost-effective diagnostic assistant, we collaborate with ED clinicians to curate a benchmark, called MIMIC-ED-Assist, that is derived from MIMIC-IV (Johnson et al., 2023b) and related datasets (Xie et al., 2022). MIMIC-ED-Assist is designed to test the ability of AI systems to provide both accurate and time-cost saving laboratory recommendations. Our benchmark consists of two prediction targets identified by our clinical collaborators to reflect patient risk: critical outcomes which includes patient death and ICU transfer (Levin et al., 2018), and lengthened ED stay, defined as ED LOS exceeding $24$ hours. Accurately identifying patients at high risks of these outcomes reduces time-cost by allowing clinicians to perform timely interventions and efficiently allocate resources. MIMIC-ED-Assist mirrors real-world ED practices by grouping individual laboratory tests into commonly performed groups, e.g., complete blood count (CBC). MIMIC-ED-Assist then tests AI systems on their ability to recommend the most informative groups to make accurate diagnostic suggestions while minimizing the total time required to perform these tests, thereby reducing LOS.
@@
With MIMIC-ED-Assist, we propose ED-Copilot which suggests a series of laboratory groups to flag patients at high risks on our prediction targets while minimizing total time-cost. ED-Copilot first linearizes (i.e., converts to text) patient information, including demographic, triage, and laboratory test results into a text sequence. It then fine-tunes a bio-medical pre-trained language model BioGPT (Luo et al., 2022) to suggest future groups and predict our two defined targets. Next, we use a reinforcement learning (RL) framework (Yu et al., 2023) to teach BioGPT to dynamically recommend the subsequent, most informative laboratory group based on prior laboratory and triage information. Unlike baselines, ED-Copilot serves as a personalized diagnostic assistant since it uses past patient information to recommend future medically relevant laboratory groups.
@@
Experiments on MIMIC-ED-Assist show that ED-Copilot outperforms state-of-the-art tree models while halving time-costs of laboratory testing from four hours to two hours. Reducing the number of laboratory tests also has the benefit of reducing financial cost. We also perform ablation studies to confirm the benefits of our feature linearization technique and the bio-medical pre-trained language model backbone. Our ablation studies also investigate the effect of size of the language model backbone on prediction accuracy which indicates larger models can lead to further gain in performance.
Our analyses also confirm the benefit of ED-Copilot’s personalized modeling approach. We show ED-Copilot can adapt its recommendations based on patient severity, thereby providing more accurate diagnostic suggestions for severely ill patients as compared to non-personalized baselines. Further, ED-Copilot achieves consistent performance across various subgroups such as age and sex. Lastly, since MIMIC-ED-Assist is a retrospective offline benchmark, we restrict ED-Copilot to only select laboratory tests a patient actually receives. We perform simulations without this restriction to approximate online performance, and show ED-Copilot is still able to make medically appropriate recommendations.
The rest of this paper is organized as follows. In Section 2, we review related work. We discuss MIMIC-ED-Assist and ED-Copilot in Sections 3 and 4 respectively. Sections 5 and 6 discuss our experimental set-up and results.
@@ 2 Related Work
Healthcare Benchmarks. Researchers have spent considerable effort in converting raw electronic health records (EHRs) into large-scale open-source datasets to ensure easy access to high-quality medical data. A notable example is the Medical Information Mart for Intensive Care (MIMIC) database (Johnson et al., 2023b) which provides patient information such as measurements, laboratory orders, and treatments, ranging from the ED to inpatient care, including the intensive care unit (ICU). MIMIC has led to the development of a range of related prediction benchmarks and models (Purushotham et al., 2018; Harutyunyan et al., 2019; Wang et al., 2020) focused on the ICU. Xie et al. (2022) took a step towards filling this gap by using the MIMIC-IV-ED (Johnson et al., 2023a) database to build a ED-focused benchmark. Their dataset includes ED triage information, and various clinical outcomes such as hospitalization that interest clinicians, and impact ED LOS.
AI Models for Healthcare. There has been significant effort to apply ML to accurately predict clinical outcomes. Traditional methods (e.g., random forests, gradient boosting, and their variants (Breiman, 2001; Chen & Guestrin, 2016; Agarwal et al., 2022, 2023)), along with deep learning (DL) have been used to predict pneumonia (Kang et al., 2020), and septic shock in the ICU (Wardi et al., 2021). Other works use interpretable models to provide diagnostic assistance in the ED such as identifying traumatic brain injury (Kornblith et al., 2022; Tan et al., 2022). Another closely related line of work is AI for cost-effective medicine. For instance, Bejnordi et al. (2017) showed DL led to faster analysis of pathology laboratory results; Komorowski et al. (2018) proposed an “AI clinician” to learn optimal dosing strategies for treating sepsis; Yu et al. (2023) focused on minimizing financial costs associated with laboratory testing while maximizing prediction accuracy. Specifically, Yu et al. (2023) used RL to sequentially select laboratory groups based on a patient’s observed test results to optimize this (financial) cost-accuracy trade-off. They validated the accuracy and cost-effectiveness of their approach on multiple clinical tasks such as diagnosing kidney injury. Researchers have also begun to explore the use of large language models (LLMs) in medical applications. LLMs have been used to extract clinical concepts (Yang et al., 2021; Luo et al., 2022; Yang et al., 2022), and facilitate medical question answering (Singhal et al., 2023; Yagnik et al., 2024).
@@ 3 MIMIC-ED-Assist Benchmark
In this section, we discuss the curation of MIMIC-ED-Assist in collaboration with ED clinicians.
@@
Task Description. We consider the following two tasks relevant to reducing ED LOS. 1) Flagging patients at high risks of critical outcome (i.e., death and ICU transfer), and ED LOS exceeding $24$ hours. 2) Providing time-cost-effective diagnostic assistance by recommending the next most medically informative laboratory group while simultaneously minimizing the time-cost of these groups.
@@
Data Pre-processing. Since laboratory results are only available for admitted patients, we filter out non-hospitalized ED patients from MIMIC-IV-ED. We only focus on adults and remove patients younger than $18$ years old. We also exclude patients that miss triage information. This step is necessary since clinicians order laboratory groups depending on triage information. To simplify the task, we remove patients who receive the same test multiple times, approximately $1.5\%$ of all patients.
@@
Clinically Relevant Outcomes. In collaboration with ED clinicians, we chose the two following prediction targets. (1) Critical outcome, which refers to death during hospitalization or transfer to an ICU within 12 hours. Identifying patients at high risks of critical outcome allows clinicians to prioritize treatment and resources for them. (2) Lengthened ED stay, indicating if ED LOS exceeding $24$ hours. Lengthened ED stay is typically correlated with the complexity of a patient’s case. Flagging patients at high risks of lengthened ED stay can enable timely intervention, and reduce ED LOS. The proportion of patients with these outcomes is described in Table 1. While these two outcomes are correlated, they also cover different aspects of patient care. For example, patients at high risks of critical outcomes often should be hospitalized quickly, while patients with complications do not necessarily have severe cases. As such, healthcare providers often require different diagnostics and resource managements for these two tasks.
@@
Triage Feature Selection. Triage features are measurements that are available for every patient before laboratory tests are ordered. We select a number of triage features in collaboration with ED clinicians. Specifically, we chose $9$ triage variables available at the beginning of patient encounters, which include patient demographics, medical history, vital signs, and chief complaints (i.e., natural language description of symptoms).
@@
Laboratory Test Selection. We only include laboratory tests performed in the ED. For simplicity, we exclude tests received by less than $5\%$ of patients and leave examination of rare tests to future research. This process results in a total of $68$ available laboratory tests. While there are $68$ tests, ED clinicians rarely order individual tests for a patient. Typically, they order groups of tests (e.g., complete blood count) based on a patient’s signs, symptoms, and risk factors. To reflect this clinical practice, our clinical collaborators categorized these 68 tests into 12 distinct groups. See Appendix A for all $68$ tests, and their assigned groupings. On average each patient receives $4.7$ laboratory groups. Consequently, MIMIC-ED-Assist contains numerous missing values, each representing a laboratory test not administered to a specific patient.
@@
ED LOS. Our benchmark records when each group is ordered, and assigns its average completion time as its ’time-cost’. ED LOS depends on the time-costs of the administered groups and can be modeled in different ways. For example, sequential tests result in ED LOS being the sum of time-costs, whereas parallel tests imply ED LOS is equal to the group with the largest time-cost. MIMIC-ED-Assist does not specify how to approximate ED LOS from these time-costs, and instead provides this flexibility to researchers and practitioners.
@@
Table 1: Statistics of MIMIC-ED-Assist. It includes information from patient triage and laboratory tests. Laboratory tests are grouped into 12 groups by ED clinicians based on how they are commonly ordered. An ED visit has a critical outcome if the patient is transferred to ICU or there is an inpatient mortality. ED stay is lengthened if the length of stay (LOS) exceeds 24 hours. The positive rate is shown in parentheses.
@@
##\## of laboratory variables 67
##\## of laboratory groups 12
Avg. ##\## of laboratory groups per patient 4.7
@@
Data Availability. Our pipeline to create MIMIC-ED-Assist from the MIMIC-IV dataset can be found at https://github.com/cxcscmu/ED-Copilot. After completing a training course and signing a data use agreement regarding patient information privacy, individuals will gain access to MIMIC-IV and can utilize our pipeline to create MIMIC-ED-Assist.
Limitations. Since MIMIC-ED-Assist is derived from public patient data, it suffers from some downsides due to a lack of data availability. (1) MIMIC-ED-Assist is derived from MIMIC-IV which only has laboratory results for hospitalized ED patients. Thus, our dataset is not reflective of the entire population that visits the ED, but instead is biased towards those with more severe issues (hospitalized). (2) As an offline benchmark derived from past patient records, all data is retrospective. As a result when developing diagnostic assistant, one can only use laboratory tests patients actually received, otherwise no testing results are available. This leads to common challenges of offline benchmarks and may result in models learning sub-optimal recommendations, and unobserved confounding. We note that measuring ‘online’ performance of ED-Copilot and AI systems in general will require clinical trials which is out of the scope of this paper. (3) MIMIC-IV is collected at the Beth Israel Deaconess Medical Center, and may not reflect the distribution of other healthcare systems. Caution is needed when generalizing models and insights derived from MIMIC-ED-Assist to new healthcare systems.